# Enhancing Zero-Shot Image Recognition in Vision-Language Models through Human-like Concept Guidance

### v1

@@ 5.2 Baselines
Table 2: Performance of zero-shot methods on ten fine-grained classification datasets. We highlight the best and second best results by bolding and underlining them, respectively.
Method SUN397 Aircraft EuroSAT Cars Food101 Pets Flower102 Caltech101 DTD UCF101 Average CLIP [6] 62.32 23.91 42.22 65.51 82.21 87.27 67.40 92.21 44.39 64.26 63.17 CLIP + E. [6] 65.13 23.67 47.73 66.27 82.28 87.11 66.10 93.06 44.04 65.00 64.13 TPT [15] 65.41 23.10 42.93 66.36 84.63 87.22 68.86 94.12 46.99 68.00 64.76 MTA [18] 64.98 25.32 38.71 68.05 84.95 88.22 68.26 94.13 45.59 68.11 64.63 Averaging Likelihood 65.54 25.74 48.96 66.51 82.46 88.14 68.41 93.91 50.65 69.13 65.95 Confidence Likelihood 66.25 26.28 48.89 66.62 82.62 90.08 71.30 93.59 51.95 69.57 66.72 TTA Likelihood 66.89 28.20 41.69 68.86 82.68 89.89 70.40 92.74 52.90 69.39 66.34
@@
Method ImageNet -A -V2 -R -Sketch Average OOD Average CLIP [6] 66.73 47.80 60.84 73.99 46.15 59.10 57.20 CLIP + E . [6] 68.38 49.95 61.91 77.71 48.26 61.23 59.46 TPT [15] 68.94 54.63 63.41 77.04 47.97 62.40 60.76 MTA [18] 69.29 57.41 63.61 76.92 48.58 63.16 61.63 Averaging Likelihood 67.42 47.84 61.00 74.59 46.37 59.44 57.45 Confidence Likelihood 67.88 48.28 61.35 74.45 46.55 59.70 57.66 TTA Likelihood 70.05 58.32 63.79 76.97 48.47 63.52 61.89
@@ 5.3 Implementation details
To obtain the concept prior, we employ GPT-4o mini to construct an importance sampler. This model is a more cost-efficient LLM than GPT-4 Turbo. We set the size of sampling windows as 4 for ten fine-grained image classification datasets ($H=4$) and 9 for the ImageNet dataset ($H=9$), aiming to enhance sampling efficiency given that ImageNet contains 1,000 classes. For all datasets, we perform 100 sampling iterations ($M_{i}=100$ for each class). The verification times for each concept in the discriminative test are set to 5 ($Z=5$). To obtain the likelihood function, the temperature parameter $\tau$ is fixed at 3 across all datasets for Confidence Likelihood. For TTA Likelihood, we follow TPT [15] to augment a single test image using random resized crops for ImageNet-series datasets and AugMix [65] for fine-grained classification datasets. This augmentation process produces a batch of 64 images ($N=64$) involving the original test image, from which we select the top 10% ($r=10$) most confident samples. Additionally, we optimize the learnable weight assigned to each concept via AdamW optimizer, using a learning rate of 1 and learning steps of 30. We adopt the default parameter settings for other baselines as reported in their respective studies. Unless otherwise specified, we use CLIP with the ViT-B/16 backbone as the primary vision-language model. To avoid randomness, we report the averaged top-1 accuracy of three different random seeds for the main experiments. The code will be made publicly available upon the acceptance of the paper.
@@ 5.5 Concept Analysis
Class Representative Concept Probability
Abbey 1 Large stone arches and a tranquil garden setting 0.2525
2 Gothic architecture and lush greenery 0.2222
3 Large stained glass windows and a tranquil garden 0.2020
4 Tall stone arches and stained glass windows 0.2020
5 Stone architecture and lush green gardens 0.1212
Outdoor Hotel 1 A lush garden and outdoor pool area 0.3000
2 Inviting outdoor seating and landscaped gardens 0.2800
3 Landscaped gardens and architectural balconies 0.1900
4 A lush garden and outdoor seating area 0.1900
5 A spacious balcony overlooking a scenic landscape 0.0400
2012 Audi A5 Coupe 1 A sleek coupe silhouette and distinctive single-frame grille 0.3200
2 A sleek, low-profile coupe design and distinctive LED headlights 0.2400
3 A sleek, aerodynamic silhouette and a prominent front grille 0.2300
4 A sleek, sporty coupe profile and distinctive Audi grille 0.1400
5 A streamlined, low-profile body and a distinctive Audi grille 0.0700
2012 Audi S5 Coupe 1 A sleek, low-profile coupe silhouette and distinctive front grille 0.3300
2 A sleek two door design prominent quad exhaust outlets 0.2800
3 A sleek coupe silhouette and aggressive sporty design 0.1600
4 A sleek, aerodynamic body and distinctive LED headlights 0.1500
5 A sporty coupe design and distinctive single-frame grille 0.0800
@@ 5.6 Ablation Study
Table 5: Performance comparison of our proposed CHBR framework on two different models (GPT-4o mini and GPT-4 Turbo) across four datasets. We highlight the best and second-best results by bolding and underlining them, respectively. Other hyperparameters are the same as in the main experiments.
@@
Model Method Pets Flower102 DTD UCF101 Average GPT-4o mini Averaging Likelihood 88.67 68.48 50.18 69.00 69.08 Confidence Likelihood 89.02 70.56 50.41 68.33 69.58 TTA Likelihood 89.51 70.52 51.65 69.13 70.56 GPT-4 Turbo Averaging Likelihood 87.02 70.44 51.83 69.23 69.63 Confidence Likelihood 88.76 71.30 52.01 69.68 70.44 TTA Likelihood 89.31 72.15 53.78 70.68 71.48
@@
Effect of size of candidate set. As discussed in Sec. 4.2, an ideal proposal distribution for concept sampling should be conditioned on both the target class $Y_{i}$ and its complement $\mathcal{Y}_{/i}$ (i.e., all classes except $Y_{i}$), formally expressed as $q(C_{i}\mid\mathcal{Y}_{/i},Y_{i})$. In practice, to approximate this distribution, we randomly sample $H$ candidate classes from the overall image class set $\mathcal{Y}$ for each target class $Y_{i}$, We then prompt LLMs to generate candidate concepts that effectively distinguish $Y_{i}$ from the selected $H$ candidate classes and the target class itself. However, this approach inherently involves a trade-off between sampling efficiency and the model’s discriminative capability. Intuitively, increasing $H$ allows for a more accurate approximation of the $q(C_{i}|\mathcal{Y}_{/i},Y_{i})$ by expanding the sampling space for concept generation while keeping sampling space unchanged. Nevertheless, a larger $H$ also necessitates the LLM to differentiate among a broader set of candidate classes, thereby increasing the complexity of the reasoning process. To assess the impact of the sampling window size $H$, we evaluate the CHBR framework using different likelihood functions across varying values of $H$ on four datasets, as illustrated in Fig. 7. The experimental results suggest that when employing the Averaging Likelihood, performance improves as $H$ increases from 4 to 9. However, beyond this point, further increases in $H$ lead to a decline in performance. Since Averaging Likelihood assumes that each concept contributes equally to the representation of the input image, it effectively reflects the quality of the sampled concept space. These findings indicate that while a larger $H$ enhances the quality of candidate concepts, an excessively large $H$ may exceed the model’s capacity, potentially leading to hallucination effects. Furthermore, in contrast to Averaging Likelihood, the other two likelihood functions exhibit greater robustness to noisy concepts, particularly when $H=19$. This observation further substantiates the effectiveness of our proposed framework, which integrates likelihood functions to dynamically adjust the weight of concepts for each input image, thereby enhancing both adaptability and reliability.
@@
Effect of size of various LLMs. Table 5 presents the performance comparison of our proposed CHBR framework on GPT-4o mini and GPT-4 Turbo [^2] https://platform.openai.com/docs/models/gpt-4-turbo, across four fine-grained classification datasets. GPT-4 Turbo may contain approximately 20 to 100 times more parameters than GPT-4o mini, as estimated based on their respective API pricing. The results indicate that GPT-4 Turbo generally outperforms GPT-4o Mini, with the highest average accuracy observed under the TTA Likelihood method. Notably, the performance gap between the models is more pronounced on datasets such as Flower102 and DTD. These findings suggest that the increased model capacity enhances the quality of the concept generation process for approximating the concept prior in real-world scenarios, ultimately leading to improved classification accuracy.
@@
Table 7: The performance comparison of our proposed CHBR framework using the original prompt and the efficient prompt across four datasets. We highlight the best and second-best results by bolding and underlining them, respectively. The sampling times for each class are set to 100 ($M_{i}=100$), and other hyperparameters are the same as in the main experiments.
Prompt Version Method Pets Flower102 DTD UCF101 Average Original Version Averaging Likelihood 88.14 68.41 50.65 69.13 69.08 Confidence Likelihood 90.08 71.30 51.95 69.57 70.73 TTA Likelihood 89.89 70.40 52.90 69.39 70.65 Efficient Version Averaging Likelihood 89.62 69.31 49.41 69.63 69.49 Confidence Likelihood 89.29 69.79 49.82 69.23 69.53 TTA Likelihood 89.86 69.63 51.30 70.13 70.23
@@ 7 Conclusion
Inspired by the human-like zero-shot image recognition process, we introduce a Concept-guided Human-like Bayesian Reasoning (CHBR) framework. CHBR models the concept used in human image recognition as latent variables and formulates this task by summing across potential concepts, weighted by a prior distribution and a likelihood. To address the trade-off between the efficiency of exploring the concept space and computational overhead, we enhance importance sampling with a discriminative test, prompting LLMs to generate concepts that aid in distinguishing candidate image classes. Additionally, we propose three likelihood functions to refine the importance of each concept, improving CHBR’s adaptability to individual test images during inference. Specially, Averaging Likelihood and Confidence Likelihood are particularly effective in fine-grained image classification, while TTA likelihood demonstrates robustness across both ImageNet and fine-grained scenarios. Nevertheless, Averaging Likelihood and Confidence Likelihood exhibit inference times comparable to directly prompting CLIP for classification, resulting in significantly lower latency compared to TTA Likelihood and other test-time adaptation baselines.

### v2

@@ 5.2 Baselines
Table 2: Performance of zero-shot methods on ten fine-grained classification datasets. We highlight the best and second best results by bolding and underlining them, respectively. The error bar of three likelihood functions is reported in Table 11 in the Appendix.
Method SUN397 Aircraft EuroSAT Cars Food101 Pets Flower102 Caltech101 DTD UCF101 Average CLIP [6] 62.32 23.91 42.22 65.51 82.21 87.27 67.40 92.21 44.39 64.26 63.17 CLIP + E. [6] 65.13 23.67 47.73 66.27 82.28 87.11 66.10 93.06 44.04 65.00 64.13 TPT [15] 65.41 23.10 42.93 66.36 84.63 87.22 68.86 94.12 46.99 68.00 64.76 MTA [18] 64.98 25.32 38.71 68.05 84.95 88.22 68.26 94.13 45.59 68.11 64.63 CHBR w/ Averaging Likelihood 65.54 25.74 48.96 66.51 82.46 88.14 68.41 93.91 50.65 69.13 65.95 CHBR w/ Confidence Likelihood 66.25 26.28 48.89 66.62 82.62 90.08 71.30 93.59 51.95 69.57 66.72 CHBR w/ TTA Likelihood 66.89 28.20 41.69 68.86 82.68 89.89 70.40 92.74 52.90 69.39 66.34
@@
Method ImageNet -A -V2 -R -Sketch Average OOD Average CLIP [6] 66.73 47.80 60.84 73.99 46.15 59.10 57.20 CLIP + E . [6] 68.38 49.95 61.91 77.71 48.26 61.23 59.46 TPT [15] 68.94 54.63 63.41 77.04 47.97 62.40 60.76 MTA [18] 69.29 57.41 63.61 76.92 48.58 63.16 61.63 CHBR w/ Averaging Likelihood 67.42 47.84 61.00 74.59 46.37 59.44 57.45 CHBR w/ Confidence Likelihood 67.88 48.28 61.35 74.45 46.55 59.70 57.66 CHBR w/ TTA Likelihood 70.05 58.32 63.79 76.97 48.47 63.52 61.89
@@ 5.3 Implementation details
To obtain the concept prior, we employ GPT-4o mini to construct an importance sampler. This model is a more cost-efficient LLM than GPT-4 Turbo [^2] https://platform.openai.com/docs/models/gpt-4-turbo. We set the size of sampling windows as 4 for ten fine-grained image classification datasets ($H=4$) and 9 for the ImageNet dataset ($H=9$), aiming to enhance sampling efficiency given that ImageNet contains 1,000 classes. For all datasets, we perform 100 sampling iterations ($M_{i}=100$ for each class). The verification times for each concept in the discriminative test are set to 5 ($Z=5$). To obtain the likelihood function, the temperature parameter $\tau$ is fixed at 3 across all datasets for Confidence Likelihood. For TTA Likelihood, we follow TPT [15] to augment a single test image using random resized crops for ImageNet-series datasets and AugMix [65] for fine-grained classification datasets. This augmentation process produces a batch of 64 images ($N=64$) involving the original test image, from which we select the top 10% ($r=10$) most confident samples. Additionally, we optimize the learnable weight assigned to each concept via AdamW optimizer, using a learning rate of 1 and learning steps of 30. We adopt the default parameter settings for other baselines as reported in their respective studies. Unless otherwise specified, we use CLIP with the ViT-B/16 backbone as the primary vision-language model. To avoid randomness, we report the averaged top-1 accuracy of three different random seeds for the main experiments. The code will be made publicly available upon the acceptance of the paper.
@@ 5.5 Concept Analysis
Class Representative Concept Probability 2012 Audi A5 Coupe 1 A sleek coupe silhouette and distinctive single-frame grille 0.3200 2 A sleek, low-profile coupe design and distinctive LED headlights 0.2400 3 A sleek, aerodynamic silhouette and a prominent front grille 0.2300 4 A sleek, sporty coupe profile and distinctive Audi grille 0.1400 5 A streamlined, low-profile body and a distinctive Audi grille 0.0700 2012 Audi S5 Coupe 1 A sleek, low-profile coupe silhouette and distinctive front grille 0.3300 2 A sleek two door design prominent quad exhaust outlets 0.2800 3 A sleek coupe silhouette and aggressive sporty design 0.1600 4 A sleek, aerodynamic body and distinctive LED headlights 0.1500 5 A sporty coupe design and distinctive single-frame grille 0.0800 Abbey 1 Large stone arches and a tranquil garden setting 0.2525 2 Gothic architecture and lush greenery 0.2222 3 Large stained glass windows and a tranquil garden 0.2020 4 Tall stone arches and stained glass windows 0.2020 5 Stone architecture and lush green gardens 0.1212 Outdoor Hotel 1 A lush garden and outdoor pool area 0.3000 2 Inviting outdoor seating and landscaped gardens 0.2800 3 Landscaped gardens and architectural balconies 0.1900 4 A lush garden and outdoor seating area 0.1900 5 A spacious balcony overlooking a scenic landscape 0.0400
@@ 5.6 Ablation Study
Table 5: Performance comparison of our proposed CHBR framework on two different models (GPT-4o mini and GPT-4 Turbo) across four datasets. We highlight the best and second-best results by bolding and underlining them, respectively. The sampling times for each class are set to 20 ($M_{i}=20$). Other hyperparameters are the same as in the main experiments.
@@
Model Likelihood Pets Flower102 DTD UCF101 Average GPT-4o mini Averaging Likelihood 88.67 68.48 50.18 69.00 69.08 Confidence Likelihood 89.02 70.56 50.41 68.33 69.58 TTA Likelihood 89.51 70.52 51.65 69.13 70.56 GPT-4 Turbo Averaging Likelihood 87.02 70.44 51.83 69.23 69.63 Confidence Likelihood 88.76 71.30 52.01 69.68 70.44 TTA Likelihood 89.31 72.15 53.78 70.68 71.48
@@
Effect of size of candidate set. As discussed in Sec. 4.2, an ideal proposal distribution for concept sampling should be conditioned on both the target class $Y_{i}$ and its complement $\mathcal{Y}_{/i}$ (i.e., all classes except $Y_{i}$), formally expressed as $q(C_{i}\mid\mathcal{Y}_{/i},Y_{i})$. In practice, to approximate this distribution, we randomly sample $H$ candidate classes from the overall image class set $\mathcal{Y}$ for each target class $Y_{i}$. We then prompt LLMs to generate candidate concepts that effectively distinguish $Y_{i}$ from the selected $H$ candidate classes and the target class itself. However, this approach inherently involves a trade-off between sampling efficiency and the model’s discriminative capability. Intuitively, increasing $H$ allows for a more accurate approximation of the $q(C_{i}|\mathcal{Y}_{/i},Y_{i})$ while keeping the sampling time unchanged. Nevertheless, a larger $H$ also necessitates the LLM to differentiate among a broader set of candidate classes, thereby increasing the complexity of the reasoning process. To assess the impact of the sampling window size $H$, we evaluate the CHBR framework using different likelihood functions across varying values of $H$ on four datasets, as illustrated in Fig. 7. The experimental results suggest that when employing the Averaging Likelihood, performance improves as $H$ increases from 4 to 9. However, beyond this point, further increases in $H$ lead to a decline in performance. Since Averaging Likelihood assumes that each concept contributes equally to the representation of the input image, it effectively reflects the quality of the sampled concept space. These findings indicate that while a larger $H$ enhances the quality of candidate concepts, an excessively large $H$ may exceed the model’s capacity, potentially leading to hallucination effects. Furthermore, in contrast to Averaging Likelihood, the other two likelihood functions exhibit greater robustness to noisy concepts, particularly when $H=19$. This observation further substantiates the effectiveness of our proposed framework, which integrates likelihood functions to dynamically adjust the weight of concepts for each input image, thereby enhancing both adaptability and reliability.
@@
Effect of size of various LLMs. Table 5 presents the performance comparison of our proposed CHBR framework on GPT-4o mini and GPT-4 Turbo, across four fine-grained classification datasets. GPT-4 Turbo may contain approximately 20 to 100 times more parameters than GPT-4o mini, as estimated based on their respective API pricing. The results indicate that GPT-4 Turbo generally outperforms GPT-4o Mini, with the highest average accuracy observed under the TTA Likelihood method. Notably, the performance gap between the models is more pronounced on datasets such as Flower102 and DTD. These findings suggest that the increased model capacity enhances the quality of the concept generation process for approximating the concept prior in real-world scenarios, ultimately leading to improved classification accuracy.
@@
Table 7: The performance comparison of our proposed CHBR framework using the original prompt and the efficient prompt across four datasets. We highlight the best and second-best results by bolding and underlining them, respectively. The hyperparameter is the same as in the main experiments.
Prompt Version Likelihood Pets Flower102 DTD UCF101 Average Original Version Averaging Likelihood 88.14 68.41 50.65 69.13 69.08 Confidence Likelihood 90.08 71.30 51.95 69.57 70.73 TTA Likelihood 89.89 70.40 52.90 69.39 70.65 Efficient Version Averaging Likelihood 89.62 69.31 49.41 69.63 69.49 Confidence Likelihood 89.29 69.79 49.82 69.23 69.53 TTA Likelihood 89.86 69.63 51.30 70.13 70.23
@@ 7 Conclusion
Inspired by the human-like zero-shot image recognition process, we introduce a Concept-guided Human-like Bayesian Reasoning (CHBR) framework. CHBR models the concept used in human image recognition as latent variables and formulates this task by summing across potential concepts, weighted by a prior distribution and a likelihood. To address the trade-off between the efficiency of exploring the concept space and computational overhead, we enhance importance sampling with a discriminative test, prompting LLMs to generate concepts that aid in distinguishing candidate image classes. Additionally, we propose three likelihood functions to refine the importance of each concept, improving CHBR’s adaptability to individual test images during inference. Specially, Averaging Likelihood and Confidence Likelihood are particularly effective in fine-grained image classification, while TTA likelihood demonstrates robustness across both ImageNet and fine-grained scenarios. Nevertheless, the inference time of Averaging Likelihood and Confidence Likelihood is comparable to that of directly prompting CLIP for classification, achieving significantly lower latency than TTA Likelihood and other test-time adaptation baselines.
@@ Appendix A Details of concept sampling algorithm
Table 11: Performance of zero-shot methods on ten fine-grained classification datasets with error bar.
Method SUN397 Aircraft EuroSAT Cars Food101 Pets Flower102 Caltech101 DTD UCF101 Average CHBR w/ Averaging Likelihood 65.54±0.02 25.74±0.07 48.96±0.34 66.51±0.11 82.46±0.02 88.14±0.05 68.41±0.24 93.91±0.17 50.65±0.29 69.13±0.18 65.95±0.02 CHBR w/ Confidence Likelihood 66.25±0.03 26.28±0.29 48.89±0.29 66.62±0.04 82.62±0.05 90.08±0.27 71.30±0.50 93.59±0.14 51.95±0.30 69.57±0.11 66.72±0.05 CHBR w/ TTA Likelihood 66.89±0.03 28.20±0.49 41.69±0.79 68.86±0.03 82.68±0.01 89.89±0.10 70.40±0.14 92.74±0.16 52.90±0.14 69.39±0.31 66.34±0.06